OpenStudy (anonymous):

Which expression is equivalent to 27x3 + 125?

OpenStudy (anonymous):

is that times 3

OpenStudy (anonymous):

156

OpenStudy (aum):

\[\large a^3 + b^3 = (a+b)(a^2-ab+b^2) \\ \large 27x^3 + 125 = (3x)^3 + 5^3 = ? \]

OpenStudy (anonymous):

just the answer?

OpenStudy (anonymous):

after the b^2) is there a plus

OpenStudy (aum):

a^3 + b^3 = (a + b)(a^2 - ab + b^2) Here, a = 3x, b = 5. Just plug them in above.